OpenStudy (natalie1234):

help please? 1. If I varies inversely with R and I = 0.25 when R = 18, which equation should be used to show this relationship? 2.xy = −16. What is the constant of variation for this relationship? 3.V varies inversely with T and V = 18 when T = 3. Which equation shows this relationship?

OpenStudy (anonymous):

solving #1, y=k/x

OpenStudy (anonymous):

so, I=k/R

OpenStudy (anonymous):

.25=k/18... solve for k

OpenStudy (natalie1234):

so its .25 = 4.5/18?

OpenStudy (anonymous):

yes! :)

OpenStudy (anonymous):

and #2, I think that the constant of variation would be the only constant in the equation... -16

OpenStudy (anonymous):

and, so you know what to do with # 3 now, right? y=k/x, V=k/T

OpenStudy (natalie1234):

yeah so for #2 thats the answer

OpenStudy (natalie1234):

i mean 1 sorry

OpenStudy (anonymous):

18=k/3, and solve for k :)

OpenStudy (anonymous):

yep :) great job! you got it :)

OpenStudy (natalie1234):

thanks!!! again :) haha

OpenStudy (anonymous):

yvw :)